Tom's age is 17 and his father's age is 50 . how many years ago was the father four times as old as tom?
Answer by ankor@dixie-net.com(12692)

Tom's age is 17 and his father's age is 50 . how many years ago was the father four times as old as tom?
:
Let x = no. years for Dad's age to be 4 times Tom's age
:
Then (17 - x) = Tom's age x years ago
:
And (50 - x) = Dad's age x years ago
:
"how many years ago was the father four times as old as tom?"
(50-x) = 4(17-x)
50 - x = 68 - 4x
+4x -x = 68 - 50
3x = 18
x = 18/3
x = 6 years ago, Dad was 4 times Tom's age
:
:
Check solution: 50 - 6 = 4(17 - 6)